OpenStudy (anonymous):

write the slope-intercept equation for the line with slope -1/2 and y-intercept (0.3) and graph the line

OpenStudy (radar):

The problem furnishes you the following; Slope m = -1/2 Point P (0,3) x=0, y=3 We have been taught the slope intercept form is: y = mx +b

OpenStudy (radar):

Couldn't you just write: y=(-1/2)x + 3 ????

OpenStudy (radar):

There was really noting to work out here, I guess the problem was to test your ability to read and comprehend.

OpenStudy (precal):

don't forget to graph the line

OpenStudy (radar):

Oh yes, there is the problem, you will need to locate another point. Find the other intercept.

OpenStudy (precal):

you can graph using the slope and y-intercept